Power Manufacturing has equipment that it purchased 5 years ago for $2,450,000. The equipment was used for a project that was intended to last for 7 years. However, due to low demand, the project is being shut down. The equipment was depreciated using the straight-line method and can be sold for $380,000 today. The company's tax rate is 34 percent. What is the after-tax salvage value of the equipment?

Multiple Choice

$509,200

$488,800

$380,000

$271,200
Answer #1

Annual depreciation=(Cost-Salvage value)/Useful Life

=(2,450,000/7)=$350,000/year

Hence book value as on date of sale=Cost-Accumulated Depreciation

=$2,450,000-(350,000*5)=$700,000

Hence loss on sale=(700,000-380,000)=$320,000

Hence after-tax salvage value=Sale proceeds+(loss on sale*Tax rate)

=380,000+(320,000*0.34)

=$488,800